OpenStudy (anonymous):

find the volume of the composite solid round your answer to the nearest tenth looks like a cylinder with half a sphere connected on each end it is 12cm long and the half shperes show 3cm halfway for radius

OpenStudy (anonymous):

Peter14?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

peachrings?

OpenStudy (anonymous):

Thanks WIO

OpenStudy (anonymous):



OpenStudy (anonymous):

It's just a sphere plus a cylinder.

OpenStudy (anonymous):

yes, can you help i have to show work and get volume

OpenStudy (anonymous):

For the cylinder use \[ \pi r^2h = \pi (3)^2(12) \]For the sphere use \[ \frac 43\pi r^3 = \frac 43\pi (3)^3 \]

OpenStudy (anonymous):

Add each part up.

OpenStudy (anonymous):

cylinder 46.1?

OpenStudy (anonymous):

sphere 123.8?

OpenStudy (anonymous):

169.9

OpenStudy (anonymous):

did you check the solution?

OpenStudy (anonymous):

I get \(144\pi \approx 452\)